A researcher predicts that watching film on institutionalization will change students'attitudes about chronically mentally ill patients. The researcher randomly selects a class of 36 students, shows them the film, and gives them a questionnaire about their attitudes. The mean score on the questionnaire for these 36 students is 70. The score for the people in general on this questionnaire is 75 with a standard deviation of 12. Using the five steps of hypothesis testing and the 5% significance level (i.e. alpha = 05), does showing the film change students'attitudes towards the chronically mentally ill?
Answer by Boreal(15235)

Ho=no change, x=75
Ha=there is a change x NE 75
alpha=0.05 P{reject Ho|Ho true}
test statistic is a t df=35,0.975=(x-75)/s/sqrt (n)
critical value |t|>2.03
calculation: t=-5*6/12
=-30/12=-2.5
This is in the rejection region, so we reject Ho and conclude at the 5% level that there was a difference.
p-value is 0.017.
The film did significantly change the student's attitudes towards the chronically mentally ill.